OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3? x-y+ 2 = 0 x+y+ 2 = 0 x+y- 2 = 0

OpenStudy (amistre64):

same equation; strip the end number x+y = ? -1 -1 = -2 x+y+2=0